OpenStudy (natalie1234):

WILL GIVE MEDALLL the first term of an arithmetic progression is -3 and the common difference is 8. what is the 28th term a. 229 b. 226 c. 220 d. 213

OpenStudy (biologicalscience):

Thank you for the medal.

OpenStudy (natalie1234):

do you know the answer?

OpenStudy (biologicalscience):

no

OpenStudy (biologicalscience):

\(x_n = a+d(n-1) \)

OpenStudy (biologicalscience):

\(\large x_{28} = -3 + 8(n-1) \rightarrow -3 +8n - 8 = 8n -11 = 213\)